A double slit produces a diffraction pattern that is a combination of single and double slit interference. Find the ratio of the width of the slits (D) to the separation between them (d), if the second minimum of the single slit pattern falls on the third-order maximum of the double slit pattern. (This will greatly reduce the intensity of the third-order maximum. Assume the central maxima of the single slit and double slit interference patterns are located at the same place on the screen.)

D/d = ?
Answer #1

For 2nd minima, destructive interference,

D sin(theta)1 = 2 lambda

For 5th maxima, constructive interference,

d sin(theta)2 = 3 lambda

Since, theta1 = theta2

D sin(theta)/d sin(theta) = 2/3

D/d = 2/3

Comment in case any doubt please rate my answer....